OpenStudy (anonymous):

(x − 2)(x + 6) = 1 Solve the quadratic equation. Give both the exact form of the answer and a calculator approximation rounded to two decimal places.

OpenStudy (anonymous):

on simplifying the equation you get, x^2 + 4x - 12 = 1 x^2 + 4x -12 -1 = 0 x^2 + 4x - 13 = 0 now you can use the formula, \[\left( \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a } \right)\] i got 2.12 and -6.12

OpenStudy (anonymous):

What were your calculations to get 2.12 and -6.12?

hero (hero):

(x-2)(x+6) = 1 x(x+6) - 2(x+6) = 1 x^2 + 6x - 2x - 12 = 1 x^2 +4x - 12 = 1 x^2 + 4x - 13 =0 (2 + d)(2 - d) = -13 4 - d^2 = -13 4 + 13 = d^2 17 = d^2 sqrt{17} = d (2 + sqrt{17})(2 - sqrt{17}) = -13 2 + sqrt{17} + 2 - sqrt{17} = 4 x^2 + 2 + sqrt{17}x + 2 - sqrt{17}x - 13 = 0 x(x + 2 + sqrt{17}) + 2-sqrt{17}(x + 2 + sqrt{17}) = 0 (x + 2 + sqrt{17})(x + 2 - sqrt{17}) = 0 x = -2 - sqrt{17} x = -2 + sqrt{-17} That's the exact solution

hero (hero):

@danieljkim12

OpenStudy (anonymous):

Very well, thank you so much!

hero (hero):

Good luck understanding the steps